OpenStudy (anonymous):

If the equation of a quadratic is y = 6x2 -4x - 1, then the equation of the axis of symmetry is:

OpenStudy (dan815):

u know what the axis of symmetry is?

OpenStudy (dan815):

dont think that sthe solution graph lol, i dunno u have to solve for it

OpenStudy (anonymous):

A. x=-1/3 B. x=1/3 C. 3 D. -3

OpenStudy (dan815):

can u find the roots of your quadratic equation!

OpenStudy (dan815):

ill show u how to get axis of syymetry form there

terenzreignz (terenzreignz):

Not the roots, @dan815 , but the vertex.

OpenStudy (dan815):

or if u want to use derivative u can get vertex straight away

OpenStudy (dan815):

other wise get the roots and find the average between them

terenzreignz (terenzreignz):

We have a vertex formula, @dan815 ... I suggest using that...

OpenStudy (anonymous):

whats the answer? a,b,c,or d?

OpenStudy (dan815):

oh theres a formula for that?

OpenStudy (dan815):

terenz there are many ways to solve this question lol, dont correct me u noob

terenzreignz (terenzreignz):

Yes, @dan815 but it is usually taken to faith, at this level... :) Can't derive that formula without calculus (specifically, derivatives, as you said) The formula is \[\Large \left(-\frac{b}{2a} \quad , \quad \frac{4ac-b^2}{4a}\right)\] And be nice.

OpenStudy (dan815):

yes but factoring :)

OpenStudy (dan815):

im just kidding btw, LOVE! <3

terenzreignz (terenzreignz):

Anyway, @micahcarr76 If your quadratic equation is of the form \[\Large y = \color{red}ax^2 + \color{blue}{b}x+\color{green}c\] Then your axis of symmetry is given by the equation \[\Large x = -\frac{\color{blue}b}{2\color{red}a}\]

OpenStudy (anonymous):

so A?

terenzreignz (terenzreignz):

I don't know. Why don't we find out? Plug the values into the formula...